OpenStudy (ineedserioushelp):

What is the equation -2x + 6y = 9 in polar form?

OpenStudy (aarontookold):

−3sinθ±3 cos2θ is that one of your answers?

OpenStudy (ineedserioushelp):

Not quite... I can tell you all the answers right now: 7√(10)/20 = r cos(Ө – 72°) 7√(10)/20 = r cos(Ө – 108°) 9√(10)/20 = r cos(Ө – 108°) 9√(10)/20 = r cos(Ө – 72°) @aarontookold

OpenStudy (iambatman):

Hey, do you know how to represent x and y using polar coordinates? \[x = r \cos \theta\]\[y = rsin \theta \] \[r^2=x^2+y^2\]